OpenStudy (moose213):

Please help will fan and medal Which property is used in the first step to solve this equation? 4 (x-6) = 5 4x - 24 = 5 4x = 29 x= 29/4 addition property of equality distributive property division property of equality commutative property

OpenStudy (moose213):

@jjcoolja12

OpenStudy (moose213):

@Qwertty123

OpenStudy (moose213):

@ganeshie8

OpenStudy (moose213):

@Ms-Brains

OpenStudy (moose213):

Please help =)

OpenStudy (qwertty123):

Hey ^.^ Do you have any guess on what to look for? If not.. you could always look up the definitions.. or I can give them.. and you can figure it out by the definition.

OpenStudy (moose213):

No...?

OpenStudy (jjcoolja12):

distributive property is my thought

OpenStudy (moose213):

What do you think @Qwertty123 ?

OpenStudy (qwertty123):

A. The property that states that if you add the same number to both sides of an equation, the sides remain equal (i.e., the equation continues to be true.) B. In general, this term refers to the distributive property of multiplication which states that the. Definition: The distributive property lets you multiply a sum by multiplying each addend separately and then add the products. C. The Division Property of Equality states that if you divide both sides of an equation by the same nonzero number, the sides remain equal. D. The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2.

OpenStudy (qwertty123):

So from that what do YOU think it is?

OpenStudy (moose213):

Bruh

OpenStudy (moose213):

distributive property...?

OpenStudy (qwertty123):

mhm ^.^

OpenStudy (moose213):

Soo yea, I'm right?

OpenStudy (jjcoolja12):

i was right too

OpenStudy (qwertty123):

Yes your right! @Moose213 and @jjcoolja12 ! :3

OpenStudy (moose213):

Thanks!

OpenStudy (qwertty123):

You Welcome ^.^

OpenStudy (jjcoolja12):

your welcome

OpenStudy (qwertty123):

Yes.. XD Nice one. :3